# Laura Barros        Aiko Schmeißer

### v1

@@ 5.1 Main results
Figure 1 reports the results from models (1) and (2) for the effect of job loss on newly registering as a party member and running for local council. We observe that both outcomes do not develop significantly differently across treated and control individuals before the layoff, supporting the parallel-trends assumption. After the layoff, outcomes sharply diverge. We estimate that job loss increases the likelihood of registering as a party member by about 0.20 percentage points, which is a 12% increase relative to the control group mean. [^23] We compare effect sizes to the mean of the control group in the post-treatment period (average across 2012, 2016, and 2020). The effect occurs in the election cycle immediately after the job loss and remains stable for the following two cycles. For the probability of running for local council, we also find a significantly positive effect that increases from 0.01 percentage points (6% relative to the control mean) in the cycle of the layoff to about 0.06 percentage points (28%) two cycles after. The dynamic effect observed for candidacies may be explained by the preparation time required for election campaigns, coupled with the fact that local councilors have no term limit and often seek re-election after their initial candidacy. In sum, the results indicate that job loss has a strong and long-lasting positive effect on individuals’ probability of engaging with politics, either as members of political parties or as local candidates.
@@
The persistence in the effect of job loss on political behavior is consistent with its effects on labor market outcomes. In Appendix Figure B.1, we show, using labor market information from RAIS, that employment and earnings effects of job loss are also long-lasting. Our estimates indicate that seven years after the layoff (the longest period we can observe with the RAIS data), dismissed workers still have a 15 percentage points lower probability of being formally employed and receive about 5,000 BRL (1,014 USD) fewer earnings per year.
@@
Robustness and external validity. Under our baseline definition of mass layoffs, there may still be considerable scope for firms to select which 30% of workers to dismiss. Appendix Table B.1 evaluates the robustness of our main findings when applying more stringent mass layoff definitions. Irrespective of whether we increase the minimum share of dismissed workers, use plant closures, increase the minimum firm size, or focus on firms with stable employment growth before the mass layoff, we still find similarly large positive effects of job loss on party memberships and candidacies. These results also mitigate concerns regarding the external validity of our analyses. The effects of mass layoffs could differ from those of regular layoffs, for example, due to spillover effects across laid-off workers or increased attention to mass layoffs by politicians or the media. The absence of differential effects by the share of displaced workers suggests that the size of a mass layoff is not an important factor driving our results. In Appendix Table B.2, we also explore the sensitivity of our results when including different sets of municipality and industry fixed effects. We show that our results are largely unaffected by accounting for local and temporal variation, suggesting that our matching strategy already compares treated and control workers in similar labor markets. Finally, for the candidacy outcome, we check the external validity of our results for individuals without a unique name. As shown in Appendix Figure B.2, the increase in the probability of running for office following job loss is very similar across individuals with and without unique names. Altogether, the robustness checks reduce concerns that our estimated effects are driven by a particular mass-layoff definition, model specification, or sample restriction.
@@
Effect timing. Next, we take a closer look at the dynamics in the estimated treatment effects. Given that deciding to become a candidate and preparing for an election campaign requires time and planning, we start by examining whether our results differ by proximity to the next local election. [^24] As mentioned in Section 2.2, in order to become candidates, individuals need to be affiliated to a political party one year prior to the election. Additionally, parties choose their candidates during party conventions, which occur a few months before the election. Appendix Figure B.3 shows results when splitting the sample by the layoff year (2009, 2010, or 2011). [^25] For this analysis and the subsequent year-level event-study model, we assign control workers the layoff year of its matched treated worker. We find that the positive effects of job loss on party memberships and candidacies are not driven by one specific treatment year, suggesting that temporal proximity to the local election does not matter. This dynamic also supports the external validity of our results. For instance, there is a concern that political responses following layoffs in 2009 could differ from other layoffs, given the proximity to the 2008 financial crisis. However, our results show similar effects for layoffs occurring in 2010 and 2011, once the crisis had already faded out in Brazil (Barbosa, 2010).
In addition, we conduct a yearly-level analysis exploiting more detailed information on the exact timing when individuals join a political party. [^26] We also rely on the year-level model when investigating the labor market consequences of layoffs (that we discussed before) as well as the consequences of layoffs for individuals’ engagement with political campaigns (which will be presented in the following paragraph). We estimate the following event-study model for new party memberships measured in each year relative to the layoff:
where $i$, $t$, and $t^{*}$ refer to workers, calendar years, and layoff years, respectively. $\delta_{k}$ capture the dynamic treatment effects for each event year after the layoff $k$ [^27] The model includes calendar year $\times$ layoff year fixed effects, $\lambda_{tt^{*}}$, to ensure that the identifying variation only comes from comparing treated to control individuals who have been matched in the same year. This stacking approach avoids the “forbidden comparisons” between late and early-treated individuals that can cause negative weights under heterogeneous treatment effects (Cengiz et al., 2019; de Chaisemartin and D’Haultfœuille, 2020; Goodman-Bacon, 2021; Sun and Abraham, 2021). Results are reported in Figure 2. We find positive and statistically significant effects on individuals’ probability of newly registering as a party member that start in the year directly following the layoff and persist up to 9 years after the shock. Reassuringly, we do not detect pre-trends before the layoff, neither in the year immediately preceding the layoff nor in earlier years. Taken together, the short-term dynamics are in line with the patterns we document at the cycle level.
@@
Third, we study the possibility of working for a political campaign. Bazzi and Labanca (2023) show that working for a winning mayoral campaign in Brazil yields substantial private income gains from improved access to jobs in the local bureaucracy. We follow the authors and identify dedicated salaried campaign staff in the RAIS data using information on the industry and legal nature of the employers. [^29] Specifically, we define campaign workers as individuals who have a full-time contract at a firm in industry Atividades de Organizações Políticas (CNAE 1.0: 91928, CNAE 2.0: 94928) or with legal nature Partido Político (3123) or Candidato a Cargo Político Eletivo (4090). Our results show an increase in individuals’ probability of working for a political campaign after job loss. Despite also being a rare event in our sample, the estimates are highly significant and suggest roughly a doubling in the likelihood of becoming a campaign worker. The effects are also highly persistent up to seven years after the layoff which is in line with our findings on party memberships and councilor candidacies.
@@ 5.2 Individual and local heterogeneity
As a next step, we analyze heterogeneity in the effects of job loss along individual characteristics. We test for differential effects by pre-layoff monthly earnings, tenure, age, education, and gender. For the first three variables, we differentiate between sample terciles, while for education, we distinguish less than 12 years (no school, elementary school, or middle school), 12 years (high school), and more than 12 years (university education). For each group, we separately estimate our DiD model (2) and scale the treatment effect by the group-specific control mean.
Results are shown in Figure 4. One striking pattern that emerges is that the effects of job loss on party memberships and candidacies are remarkably pervasive, being significantly positive in all the individual sub-samples considered. Appendix Figure B.8 shows that the effects are also pervasive across several municipality characteristics. These include the local average household income, unemployment rate, Gini index, homicide rate, as well as the electorate’s ideological leaning, which is measured as the vote shares for the left- vs. center-right candidate (Luiz Inácio Lula da Silva vs. Geraldo Alckmin) in the last pre-treatment Presidential election of 2006. Despite the large heterogeneity across regions in Brazil, we find significant positive effects of job loss on political entry along the distribution of all these local characteristics.
@@
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) for different subgroups of individuals. For education, we differentiate no degree, elementary school, or middle school ($<$ 12 years), high school (12 years), and university education ($>$ 12 years). For age, pre-layoff monthly wage, and tenure, the sample is split by terciles of each variable. The horizontal bars depict 95% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the group-specific mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
@@
Beyond the overall positive effects, Figure 4 reveals substantial differences across educational groups with higher effects on party memberships and candidacies for better-educated individuals. In quantitative terms, we estimate that individuals with university education experience an increase of about 25% in both the probability of joining a political party and running for office, while for individuals with lower schooling levels, the effects on both outcomes are smaller than 10%. These patterns are consistent with the opportunity-cost framework presented before. Highly educated individuals have better earnings prospects in the private sector. Being displaced due to a mass layoff may thus imply a larger shock to their opportunity costs in terms of increased time availability and reduced labor income. Furthermore, benefits from political engagement may be higher for individuals with higher levels of education, making a political career more attractive. These results may also have important implications for the competence of individuals who choose to enter politics. We will examine the consequences of layoffs for political selection in more detail in Section 7.
@@
For the effects on new party memberships, we do not observe a clear gradient for any of the other individual characteristics considered. In contrast, for the impact on candidacies, we find higher effects for younger and low-tenured individuals and for women. The larger increases for young workers are consistent with the “impressionable years” hypothesis according to which political attitudes are more malleable at younger ages (e.g., Krosnick and Alwin, 1989). Moreover, the labor market careers of younger individuals are less entrenched, making it easier to start a political position after the layoff. Given the large gender gaps in labor market opportunities, careers are also less entrenched for women. This potentially explains women’s higher likelihood of transitioning to a political career after a layoff.
@@ 5.3 Monetary incentives
Holding a political position often brings monetary benefits. Running for political office may thus be an appealing strategy for mitigating the income loss resulting from being laid off in the private sector. To examine the role of financial incentives in shaping the decision to enter politics after a job loss, we conduct two heterogeneity analyses. First, we study the attractiveness of local councilor jobs by exploiting variation in councilor wages across Brazilian municipalities. Second, to capture changes in private-sector opportunity costs, we construct a measure of the predicted income loss that arises from job loss and test whether it relates to differences in job loss effects on political participation.
@@
In Figure 5, we examine heterogeneity in the effects of job loss by the maximum wage that councilors can receive in a municipality. We find a strong gradient in the effects on running for councilor. In municipalities with the lowest wage caps, job loss does not significantly affect local candidacies. In contrast, in municipalities in which councilors can earn up to 75% of state deputy salaries, job loss is estimated to increase the likelihood of running for councilor by more than 30% relative to the control mean. At the same time, we do not find similar patterns for the effects on party memberships. As party members do not receive wages, this result is expected and gives us confidence that we are capturing the role of financial incentives, rather than other differences between small and large municipalities, in influencing the decision to enter politics after being laid off.
@@
Predicted income loss. Next, we examine whether predicted income losses due to the layoff are related to the effects of job loss on political entry. [^30] We focus on predicted, rather than actual earnings loss, since the latter may be directly affected by individuals’ behavior after the layoff, such as their decision to run for office. Hilger (2016) and Britto et al. (2022a) conduct similar exercises for the effect of parental job loss on children’s outcomes. The income effects of job loss are predicted as follows. Similar to Schmieder et al. (2023), we first construct for each laid-off worker in our sample an individual-specific measure of the relative earnings losses after layoff by exploiting the fact that we have matched each treated worker to a similar control worker:
where $\bar{y}_{i,post}$ is the average earnings in the 5 years after the layoff of treated worker $i$, or her/his matched control worker $i^{\prime}$. $\bar{y}_{i,pre}$ is the average earnings in the 5 years before the layoff. This double difference can be thought of as the share of earnings that each worker loses in the medium run because of the layoff. We then regress $\Delta_{dd}y_{it}$ on a rich set of pre-layoff characteristics and compute the predicted relative earnings loss. The regression includes the following individual covariates: gender, education, age decile, earnings decile, tenure decile, firm size decile, and layoff year. We also add municipality $\times$ industry dummies to capture variation in labor market conditions. Finally, we re-estimate the effects of job loss on new party memberships and candidacies in model (2), by splitting the sample into quintiles of the predicted earnings loss distribution.
Figure 6: Heterogeneity by predicted earnings loss
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B). We split the sample by quintiles of the treated individuals’ predicted relative earnings loss due to the layoff. See the text for details on how earnings losses are predicted. Control workers are assigned to their matched treated workers in each sub-sample. The horizontal bars depict 95% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the group-specific mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
Results are shown in Figure 6. The predicted income losses are clearly correlated with the estimated effects of job loss on councilor candidacies. In the lowest quintile of earnings losses, we find null effects, whereas in the highest quintile job loss is estimated to raise the probability of running for councilor by more than 30%. Again, we do not find heterogeneity in the effects on new party memberships. These results offer further support to the idea that the reduction in private-sector opportunity costs is an important mechanism in explaining the positive effect of job loss on pursuing a political office that comes with financial benefits.

### v3

@@ 4.2 Main results
We begin by discussing the effects of job loss on labor market trajectories. Figure 1 displays yearly event-study estimates based on model (3). In the year of the layoff, workers experience a sharp decline in labor earnings of around 10,000 BRL (2,028 USD), which correspondents to a 45% reduction relative to baseline earnings in the year prior to the layoff. Moreover, the probability to be formally employed at the end of the year falls by about 60 percentage points. Starting two years after the layoff, both earnings and employment start to continuously recover; yet, the recovery is only partial and does not return to pre-layoff levels. Seven years after the layoff – the longest horizon we can observe in the RAIS data – dismissed workers still earn about 5,000 BRL (1,014 USD) less per year and are 15 percentage points less likely to be formally employed compared to the control group. Appendix Figure B.1 further shows that job loss reduces subsequent job stability, as reflected in a higher frequency of job separations. These findings align with the job loss literature that has widely documented the long-term scarring effects of job displacements (e.g., Jacobson et al., 1993; Eliason and Storrie, 2006; Davis and von Wachter, 2011; Lachowska et al., 2020; Schmieder et al., 2023). [^26] Britto et al. (2022b) also show, using survey data on Brazilian workers, that accounting for income from the informal sector reduces the estimated earnings losses following layoffs by only about 10%.
Figure 1: Effect of job loss on labor earnings and employment
A. Labor earnings B. Employment Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on yearly formal labor earnings (Panel A) and the likelihood of being formally employed at the end of the year (Panel B). $N=24,549,564$ individual-year observations. The graphs include vertical bars that depict 95% confidence intervals based on standard errors clustered at the individual level (too small to be visible). Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group across all post-treatment periods are reported.
@@
Figure 2 presents our main results on the effects of job loss on newly registering as a party member and running for local council, based on estimates from models (1) and (2). We observe that both outcomes do not develop significantly differently across treated and control individuals before the layoff, lending support to the parallel-trends assumption. After the layoff, outcomes sharply diverge. We estimate that job loss increases the likelihood of registering as a party member by about 0.20 percentage points, which is a 12% increase relative to the control group mean. [^27] We compare effect sizes to the mean of the control group in the post-treatment period (average across 2012, 2016, and 2020). The effect emerges in the election cycle immediately after the job loss and remains stable in the subsequent two cycles. For the probability of running for local council, we also find a significantly positive effect that increases from 0.01 percentage points (6% relative to the control mean) in the cycle of the layoff to about 0.06 percentage points (28%) two cycles later. The gradual increase in effects on candidacies may reflect the preparation time required for election campaigns, coupled with the fact that local councilors have no term limit and often seek re-election after their initial candidacy. Moreover, the persistence of effects over three election cycles (12 years) is consistent with the long-term decline in private-sector employment prospects documented in Figure 1, which may increase the relative attractiveness of political careers. In sum, the results indicate that job loss has a strong and long-lasting positive effect on individuals’ probability of engaging with politics, either as members of political parties or as local candidates.
@@
Robustness and external validity. Under our baseline definition of mass layoffs, there may still be considerable scope for firms to select which 30% of workers to dismiss. Appendix Table B.2 evaluates the robustness of our main findings when applying more stringent mass layoff definitions. Irrespective of whether we increase the minimum share of dismissed workers, use plant closures, increase the minimum firm size, or focus on firms with stable employment growth before the mass layoff, we still find similarly large positive effects of job loss on party memberships and candidacies. These results also mitigate concerns regarding the external validity of our analyses. The effects of mass layoffs could differ from those of regular layoffs, for example, due to spillover effects across laid-off workers or increased attention to mass layoffs by politicians or the media. The absence of differential effects by the share of displaced workers suggests that the size of a mass layoff is not an important factor driving our results. In Appendix Table B.3, we also explore the sensitivity when including different sets of municipality and industry fixed effects. We show that our results are largely unaffected by accounting for local and temporal variation, suggesting that our matching strategy already compares treated and control workers in similar labor markets. Finally, for the candidacy outcome, we check the external validity of our results for individuals without a unique name. As shown in Appendix Figure B.2, the increase in the probability of running for office following job loss is very similar across individuals with and without unique names. Altogether, the robustness checks reduce concerns that our estimated effects are driven by a particular mass-layoff definition, model specification, or sample restriction.
@@
Effect timing. Next, we take a closer look at the dynamics in the estimated treatment effects. Given that deciding to become a candidate and preparing for an election campaign requires time and planning, we start by examining whether our results differ by proximity to the next local election. Appendix Figure B.3 shows results when splitting the sample by the layoff year (2009, 2010, or 2011). We find that the positive effects of job loss on party memberships and candidacies are not driven by one specific treatment year, suggesting that temporal proximity to the local election does not matter. This dynamic also supports the external validity of our results. For instance, there may be the concern that political responses following layoffs in 2009 could differ from other layoffs, given the proximity to the 2008 financial crisis. However, our results show similar effects for layoffs occurring in 2010 and 2011, once the crisis had already faded out in Brazil (Barbosa, 2010).
In addition, we conduct a year-level analysis exploiting more detailed information on the exact timing when individuals join a political party. Figure 3 presents results from the yearly event-study estimated using model (3). We find positive and statistically significant effects on individuals’ probability of newly registering as a party member that start in the year directly following the layoff and persist up to 9 years after the shock. Reassuringly, we detect no evidence of pre-trends prior to the layoff – neither in the year immediately preceding the layoff nor in earlier years. Taken together, the short-term dynamics are consistent with the patterns observed at the cycle level.
@@
Third, we study the possibility of working for a political campaign. Bazzi and Labanca (2023) show that working for a winning mayoral campaign in Brazil yields substantial private income gains from improved access to jobs in the local bureaucracy. Following their approach, we identify dedicated salaried campaign staff in the RAIS data using information on the industry and legal nature of the employers. [^29] Specifically, we define campaign workers as individuals who have a full-time contract at a firm in industry Atividades de Organizações Políticas (CNAE 1.0: 91928, CNAE 2.0: 94928) or with legal nature Partido Político (3123) or Candidato a Cargo Político Eletivo (4090). Our results show an increase in individuals’ probability of working for a political campaign after job loss. Despite also being a rare event in our sample, the estimates are highly significant and suggest roughly a doubling in the likelihood of becoming a campaign worker. The effects are also highly persistent up to seven years after the layoff which is in line with our findings on party memberships and councilor candidacies.
@@ 4.3 Individual and local heterogeneity
As a next step, we analyze heterogeneity in the effects of job loss along key socio-demographic characteristics. Specifically, we test for differential effects by pre-layoff education (categorized into three groups: middle school or less, high school, university education), age (in terciles), and gender. For each subgroup, we separately estimate the DiD model from equation (2) and scale the treatment effect by the group-specific mean of the control group.
Results are shown in Figure 5. One striking pattern that emerges is that the effects of job loss on party memberships and candidacies are remarkably pervasive, being significantly positive in all individual subsamples considered. Appendix Figure B.4 shows that the effects are also pervasive across several municipality characteristics. These include the local average household income, unemployment rate, Gini index, homicide rate, as well as the electorate’s ideological leaning, which is measured as the vote shares for the left- vs. center-right candidate (Luiz Inácio Lula da Silva vs. Geraldo Alckmin) in the last pre-treatment Presidential election of 2006. Despite the large heterogeneity across regions in Brazil, we find significant positive effects of job loss on political entry along the distribution of all these local characteristics.
@@
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) for different subgroups of individuals. For education, we differentiate no degree, elementary school, or middle school ($<$ 12 years), high school (12 years), and university education ($>$ 12 years). For age, the sample is split by terciles. The horizontal bars depict 95% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the group-specific mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
@@
Beyond the overall positive effects, Figure 5 reveals substantial differences across education levels, with larger effects on party membership and candidacy among more educated individuals. In quantitative terms, we estimate that individuals with a university degree ($>$ 12 years of education) experience an increase of about 25% in the probabilities of both joining a political party and running for office, while for individuals without a high school degree ($<$ 12 years of education), the estimated effects on both outcomes are below 10%. These patterns are consistent with the opportunity-cost framework presented earlier. Highly educated individuals tend to have better earnings prospects in the private sector, so being displaced in a mass layoff may represent a larger shock to their labor income. Furthermore, the benefits from political engagement – whether in terms of prestige, perceived chances of winning, or future job opportunities – may be higher for more educated individuals, making a political career more attractive. To shed light on these mechanisms, Appendix Figure B.5 examines post-layoff earnings and employment trajectories by education level. The results show that highly-educated individuals indeed suffer substantially larger total earnings losses after the layoff. While their overall employment rate does not improve more rapidly, they are more likely to transition into public-sector jobs, suggesting that political connections may offer them greater potential returns through improved access to such positions.
@@
In terms of the other two individual characteristics considered – age and gender – we do not observe a clear gradient in the effects on new party memberships. In contrast, the impact on candidacies is more pronounced among younger individuals and women. These patterns may reflect differences in labor market attachment and career flexibility. Younger workers often face less entrenched career paths, which can lower the barriers to start a political position after a layoff. [^30] The larger effects for young workers are also consistent with the “impressionable years” hypothesis according to which political attitudes are more malleable at younger ages (e.g., Krosnick and Alwin, 1989). Similarly, given persistent gender gaps in labor market opportunities, women may also face fewer opportunity costs when shifting toward a political career, contributing to their higher likelihood of running for office following displacement.
@@ 4.4 Monetary incentives
Holding a political position often entails substantial monetary rewards, making it an appealing strategy for mitigating the income loss resulting from being laid off in the private sector. To examine the role of financial incentives in shaping the decision to enter politics after a job loss, we conduct two heterogeneity analyses. First, to capture changes in private-sector opportunity costs, we construct a measure of the predicted income loss from displacement and test whether individuals facing larger losses are more likely to enter politics. Second, we study the attractiveness of local councilor jobs by exploiting variation in councilor wages across Brazilian municipalities.
Predicted income loss. We start by examining whether predicted income losses due to the layoff are related to the effects of job loss on political entry. [^31] We focus on predicted, rather than actual earnings loss, since the latter may be directly affected by individuals’ behavior after the layoff, such as their decision to run for office. Hilger (2016) and Britto et al. (2022a) conduct similar exercises for the effect of parental job loss on children’s outcomes. The income effects of job loss are predicted as follows. Similar to Schmieder et al. (2023), we first construct for each laid-off worker in our sample an individual-specific measure of the relative earnings losses after layoff by exploiting the fact that we have matched each treated worker to a similar control worker:
where $\bar{y}_{i,post}$ is the average earnings in the 5 years after the layoff of treated worker $i$, or her matched control worker $i^{\prime}$. $\bar{y}_{i,pre}$ is the average earnings in the 5 years before the layoff. This double difference can be thought of as the share of earnings that each worker loses in the medium run because of the layoff. We then regress $\Delta_{dd}y_{it}$ on a rich set of pre-layoff characteristics and compute the predicted relative earnings loss. The regression includes the following individual covariates: gender, education, age decile, earnings decile, tenure decile, firm size decile, and layoff year. We also add municipality $\times$ industry dummies to capture variation in labor market conditions. Finally, we re-estimate the effects of job loss on new party memberships and candidacies in model (2), by splitting the sample into quintiles of the predicted earnings loss distribution.
Figure 6: Heterogeneity by predicted earnings loss
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B). We split the sample by quintiles of the treated individuals’ predicted relative earnings loss due to the layoff. See the text for details on how earnings losses are predicted. Control workers are assigned to their matched treated workers in each subsample. The horizontal bars depict 95% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the group-specific mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
Results are shown in Figure 6. The predicted income losses are clearly related to the estimated effects of job loss on councilor candidacies. In the lowest quintile of earnings losses, we find null effects, whereas in the highest quintile job loss is estimated to raise the probability of running for councilor by more than 30%. In contrast, we do not find heterogeneity in the effects on new party memberships. These results support the idea that the reduction in private-sector opportunity costs is a key mechanism behind the rise in political entry after layoffs – particularly for time-intensive forms of political engagement such as running for office, where the opportunity cost of foregone earnings plays a larger role than for lower-commitment activities like party membership.
@@
In Figure 7, we examine heterogeneity in the effects of job loss by the maximum wage that councilors can receive in a municipality. We find a strong gradient in the effects on running for councilor. In municipalities with the lowest wage caps, job loss does not significantly affect local candidacies. In contrast, in municipalities in which councilors can earn up to 75% of state deputy salaries, job loss is estimated to increase the likelihood of running for councilor by more than 30% relative to the control mean. Again, we do not find similar patterns for the effects on party memberships. As party members do not receive wages, this result is expected and gives us confidence that we are capturing the role of financial incentives, rather than other differences between small and large municipalities, in influencing the decision to enter politics after being laid off.